OpenStudy (anonymous):

Can someone explain to me how to do this problem please? What is the equation of the given linein standard form? Use interger coefficents. y=-1.7x+8.5

OpenStudy (anonymous):

Hint: It's in Slope-Intercept form: y = mx + b You need to change it to standard form: ax + by = c So: y= -1.7x + 8.5 + 1.7x 1.7x + y = 8.5

OpenStudy (anonymous):

my options are- 17x+10y=85 -17x+10y=85 -17x+10y=-85 17x+10y=-85

OpenStudy (anonymous):

@TuringTest @hartnn

OpenStudy (anonymous):

??? whats that?

OpenStudy (anonymous):

Sorry, I blanked out on how to do this.

hartnn (hartnn):

u have 1.7x + y = 8.5 right ? now multiply 10 on both sides.,what u get ?

OpenStudy (anonymous):

ahhh its alright.. lol 17x+10y=85?

hartnn (hartnn):

thats correct :)

OpenStudy (anonymous):

thank you! I have one more question i need help with

hartnn (hartnn):

post it as a new question.